OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1. Show your work for full credit.

OpenStudy (anonymous):

\[\theta = - \frac{ 3\pi }{ 4 }\]

OpenStudy (anonymous):

@ganeshie8 @satellite73

ganeshie8 (ganeshie8):

can you plot the terminal side that represents the given angle, \(-\dfrac{3\pi}{4}\) ?